OpenStudy (anonymous):

How to use a table of values to solve a system of equation? y = x - 1 and 2x + y = 11 ? WILL MEDAL AND FAN ;)

OpenStudy (anonymous):

@jeweaver6895

OpenStudy (anonymous):

do you need to solve for a specific value of x and y?

OpenStudy (anonymous):

Well, under the x in the table are the numbers 1,2,3,4,5. the y is blank.

OpenStudy (anonymous):

oh ok, so what you want to do is take the values of x that you are given and then put them into your equation y=x-1 for x and solve for y for each value of x

OpenStudy (anonymous):

Oh could you maybe walk through it with me?

OpenStudy (anonymous):

What about the 2x + y = 11 equation btw?

OpenStudy (anonymous):

ok lets start with x=1, use the equation y=x-1 and substitute 1 in for x and solve for y to get y=(1)-1, or y=0, repeat that for the x values to fill in the chart. From the information given, im not sure what to do with the second equation

OpenStudy (anonymous):

Maybe we first solve both the equations and then use that to finish the table?